| Operation Summary for December 16, 2011 - December 19, 2011 |
|
NuMI reported receiving 9.11E18 protons on target for the period from 12/12/11 to 12/19/11. Friday December 16 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, and with the MiniBooNE and NuMI experiments taking beam. During the day shift, At 8:41 AM, Operators rebooted a NuMI BPM VME (NBPM6N). At 9:44 AM, the MiniBooNE beam permit tripped off for the first of two times in about an hour and a half due to bad beam positions; it reset. A Linac Klystron RF station (KRF2) tripped off for the first of three times in 11 minutes; it reset. At 10:03 AM, Operators reported that a Linac quadrupole power supply’s output (L:QPS107 ) was wandering; a Linac person began repairs. Muon Ring personnel began a study at 10:40 AM. Operations established beam to Meson FTBF experiment T-1015 at 11:16 AM. A Linac technician completed his work on L:QPS107 at 12:31 PM. The Muon Ring study ended at 2:33 PM. At 2:58 PM, the NuMI beam permit tripped off due to an incomplete MI RF gap; it reset. During the evening shift, at 4:55 PM, KRF2 tripped off; it reset. At 6 PM, LRF7, which was the test stand, was causing problems for KRF2; a Linac person adjusted LRF7. LRF2 tripped off at 9:25 PM; it reset. Operations disable T-1015 beam at 10:05 PM. Saturday December 17 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 2:33 AM, the MiniBooNE beam permit tripped off for the first of four times in about two and half hours a due to bad beam positions; it reset. KRF1 tripped off at 4:29 AM; it reset. At 7:53 AM, Operators reported that the H- Source output (the backup source) began falling; Operators increased the source pressure and the arc supply voltage. At 7:56 AM, Operators reported that the output of the I- Source had dropped twice in the last four hours; operators tuned the extractor and magnet to bring the output up. During the day shift, Operations established beam to T-1015 at 10:27 AM. At 3:31 PM, the NuMI beam permit tripped off due to an incomplete MI RF gap; it reset. During the evening shift, LRF4 tripped off at 4:12 PM; it reset. At 9:28 PM, LRF1 tripped off on a driver ready fault; Operators investigated, found a LCW leak, and called in Linac personnel. Operations disable T-1015 beam at 10:23 PM. Sunday December 18 The midnight shift began with beam off due to LRF1 repairs. During the midnight shift, Operators reported that LRF1 had been repaired and Operations established beam to MiniBooNE and NuMI at 6:31 AM. During the day shift, KRF3 tripped off at 9:53 AM; it reset. Operations established beam to T-1015 at noon. At 2:58 PM, the NuMI beam permit tripped off due to an incomplete MI RF gap; it reset. During the evening shift, KRF1 tripped off at 5:57 PM; it reset. MIRF9 tripped off at 6:39 PM; it reset. Operations disable beam to T-1015 shortly before midnight. Monday December 19 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 3:40 AM, the MiniBooNE beam permit tripped off for the first of two times in about 45 minutes due to bad beam positions; it reset. At 6:35 AM, LRF4 tripped off for the first of three times in about an hour; it reset. Operations Coordinator (OpsCo) The OpsCo approved of the plans by Booster and Muon Ring personnel to conduct studies. Beam will be shut off at 7 AM on Tuesday, 12/20/11, for a six-hour access for maintenance, necessary repairs, and Interlock testing. |
|
More Information Comments and Suggestions |
| last modified 12/19/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |